If a certain number is added to the numerator and denominator of 715, the result is 1/3, Find the number?
Answer by Alan3354(69443)

(x+715)/(x+1) = 1/3
x+1 = 3x + 2145
2x = -2144
x = -1072
---------
(715-1072)/-1071 = -357/-1071 = 1/3